# Experimental validation of a linear momentum and bluff-body model for high-blockage cross-flow turbine arrays

### v1

@@ 5 Prediction of Confined Performance
To quantify the effectiveness of this forecasting method, the measured array performance at $\beta_{1}=30.0\%-55.0\%$ (figure 6) is used to predict array performance at $\beta_{2}$ where measurements are also available. The resulting $C_{T,2}$ and $C_{P,2}$ at each operating point are then compared to the experimentally measured performance (denoted as $C_{T,0}$ and $C_{P,0}$) at each $\beta_{2}$. The predicted thrust coefficients at $\beta_{2}=30.0\%$, $45.0\%$ and $55.0\%$ and the corresponding difference between the predicted and measured thrust coefficients are given in figure 19. The measured thrust coefficients are predicted within $\sim\!0.2$ at $\beta_{2}=30.0\%$ and $45.0\%$ (figures 19 and 19) and within $\sim\!0.3$ at $\beta_{2}=55.0\%$ (figure 19). In general, the magnitude of the thrust prediction error increases with the difference between $\beta_{1}$ and $\beta_{2}$ (i.e., predictions further from the starting blockage are less accurate), and the absolute prediction error tends to grow as $\beta_{2}$ increases for the same difference between $\beta_{1}$ and $\beta_{2}$. In relative terms, these errors are appreciable at lower $\lambda_{2}$ where $C_{T}$ is relatively low, but decline for higher $\lambda_{2}$. Since the foundation of this forecasting method is Maskell-inspired bluff-body scaling (Eqs. 16, 17 and 18), the differences in the predicted $C_{T,2}$ across $\beta_{1}$ are driven by differences in ${C_{T}}^{b}$ across $\beta_{1}$. Thus, the shapes and relative positions of the ${C_{T}}^{b}-\lambda^{b}$ curves across $\beta_{1}$ (figure 17) dictate the shapes and relative positions of the resulting $C_{T,2}-\lambda_{2}$ curves when scaled to $\beta_{2}$ (figures 19, 19 and 19). For the current data set, at low to moderate tip-speed ratios this corresponds to underpredicted $C_{T,2}$ if $\beta_{1}>\beta_{2}$ and overpredicted $C_{P,2}$ if $\beta_{1}<\beta_{2}$, except at relatively high $\lambda$ when this trend inverts.
@@
The corresponding predicted power coefficients and prediction error are shown in figure 20. The measured power coefficients are predicted within $\sim\!0.05$ at $\beta_{2}=30.0\%$ (figure 20), $\sim\!0.1$ at $45.0\%$ (figure 20), and $\sim\!0.2$ at $\beta_{2}=55.0\%$ (figure 20). The greatest prediction error tends to occur near the performance peak and at tip-speed ratios well beyond the optimal operating point. However, the tip-speed ratio corresponding to maximum performance is well-predicted (figures 20, 20 and 20). Since only the thrust influences the outcome of the bluff-body forecasting procedure, trends in the $C_{P,2}-\lambda_{2}$ curves with $\beta_{1}$ resemble those for the ${C_{T}}^{b}-\lambda^{b}$ curves.
@@
The proposed forecasting method combines linear momentum theory and the bluff-body theory of Maskell to allow for analytically predicting confined turbine performance from experimental or simulation data. However, since this method utilizes $u_{b}$ as a link between $\beta_{1}$ and $\beta_{2}$, the prediction accuracy depends on whether the bypass velocity is descriptive of array performance at both $\beta_{1}$ and $\beta_{2}$. Consequently, the effectiveness of this forecasting scheme outside of the tested range of $\beta$ is unknown, and the method is likely less accurate when turbine dynamics depart from a bluff-body model. If there exists a transitional condition at which both a Maskell-inspired LMADT scaling (figure 17) and traditional Glauert-derived LMADT scaling (figure 17) provide similar results, then the use of a hybrid method that translates between these flow regimes might allow for forecasting performance across a wider ranges of blockages. However, the existence of such a transitional condition is speculative.
6 Conclusions
@@
This research investigates the performance and near-wake flow fields of a cross-flow turbine array at the upper end of practically achievable blockage ratios and demonstrates the potential of theoretical frameworks to describe the dynamics of real turbines under these conditions. The performance and near-wake flow field of a two-turbine array were experimentally characterized over a wide range of operating conditions at array blockage ratios from $30\%$ to $55\%$, holding other non-dimensional flow parameters constant to isolate blockage effects. The observed array performance supports the trends observed in prior work while providing a uniquely high-resolution view of the manner in which performance changes with the blockage ratio at relatively high confinements. This highlights relatively subtle trends, such as linear behavior of both the power and thrust coefficients with blockage for low-to-moderate $\beta$ and $\lambda$. Similarly, the measured near-wake velocity fields quantitatively illustrate how the wake and bypass flows change with the blockage and tip-speed ratios. Critically, this work provides an openly-available set of experimental performance and flow-field data for cross-flow turbines at high blockage, which will aid in the validation of numerical simulations and reduced-order models that seek to examine confinement effects and inform the design of full scale systems that can exploit these physics.
@@
A key outcome of this work is the demonstration that linear momentum actuator disk models can quantitatively describe aspects of real turbine arrays in highly confined flows. It is shown that, at a location $0.6$ turbine diameters downstream of the array, the velocity of the flow bypassing the array measured using PIV is well represented by the bypass velocity predicted from the open-channel LMADT model of Houlsby et al. (2008) with the measured array thrust coefficients and flow conditions as inputs. This result was confirmed for a wide range of blockages and tip-speed ratios, and the bypass velocity calculated from LMADT falls within approximately $8\%$ of the measured velocity between the turbines at the array centerline at all conditions. The residual deviation between modeled and measured bypass velocities may be a consequence of near-wake mixing that is neglected in LMADT. Following Whelan et al. (2009) and inspired by the bluff-body theory of Maskell (1963), scaling the array power and thrust coefficients by this analytical bypass velocity collapses these coefficients across the tested blockage and tip-speed ratios. These results highlight similar dynamics between high-blockage turbine arrays and bluff bodies, and demonstrate that, despite their inherent simplicity, LMADT models can effectively model aspects of the flow around non-ideal turbines in confined flows and provide insights into the salient hydrodynamics that govern their performance. Leveraging the connection between array performance and the bypass velocity, an analytical forecasting method based on LMADT is presented by which known turbine performance at one confinement ($\beta_{1}$) can be used to predict turbine performance at a different confinement ($\beta_{2}$) with moderate accuracy. The presented forecasting method provides a pathway for estimating turbine performance under confined conditions based on the dominant dynamics of high-blockage arrays. As such, we expect the broader applicability of this method and accuracy as a blockage correction to be influenced by factors such as rotor geometry and rotation rate that influence the similarity between turbine and bluff body dynamics (Araya et al., 2017).
@@
Although this work focuses on cross-flow turbines, the characteristic fluid dynamics of confinement are not exclusive to this archetype, and the analytical models employed are agnostic to turbine geometry. However, the design of the rotor is expected to influence how these characteristic dynamics change with blockage and the tip-speed ratio (Schluntz & Willden, 2015), such that the effectiveness of Maskell-inspired bluff-body scaling and forecasting likely depend on turbine geometry and operating condition. It is recommended that future work explore the broader application of these analytical methods to turbines at lower confinement, as well as to different turbine designs. Additionally, although this work focuses on the time-averaged array performance and near-wake flow fields, the unsteady fluid dynamics of cross-flow turbines also inform their performance and the properties of the near-wake. Future research should examine how the phase-resolved performance of cross-flow turbines develops with confinement, and how these dynamics influence the temporal and spatial evolution of the flow field in the near- and far-wake regions.
[…]
@@
[Acknowledgements]The authors would like to thank Gregory Talpey, Gemma Calandra, and Corey Crisp for their assistance in commissioning the high-confinement test-rig, as well as help with data collection. Abigale Snortland, Hannah Ross, Jennifer Franck, and Mukul Dave are acknowledged for insightful conversations regarding the results in this work. We are grateful to the Alice C. Tyler Charitable Trust for upgrades to the experimental flume, including closed loop heating and cooling.
[…]
[Funding]This work was supported by the United States Advanced Research Projects Agency – Energy (ARPA-E) under award number DE-AR0001441.
[…]
[Declaration of interests]The authors report no conflict of interest.
[…]
@@
[Data availability statement] The data that support the findings of this study are openly available in the Dryad data repository at https://doi.org/10.5061/dryad.dfn2z35b5 (DOI pending activation). The MATLAB code used to implement open-channel linear momentum theory is available on GitHub at https://github.com/aidan-hunt/linear-momentum.
[…]
@@
[Supplementary data] The following are provided as supplementary material: lateral force measurements, estimates of hydraulic efficiency, additional visualizations of the flow fields at $0.6D$ and $1.5D$ downstream of the array, and a visualization of the solution space for the bluff-body forecasting method.
[…]
@@
[Author contributions (CRediT)] A. Hunt: Conceptualization, Investigation, Validation, Methodology, Formal Analysis, Visualization, Software, Writing - Original Draft and Review & Editing. A. Athair: Investigation, Validation, Methodology, Formal Analysis, Visualization, Software, Writing - Original Draft and Review & Editing. O. Williams: Conceptualization, Methodology, Supervision, Writing - Review & Editing B. Polagye: Conceptualization, Methodololgy, Supervision, Writing - Original Draft and Review & Editing, Resources, Funding Acquisition
[…]
[Author ORCIDs]A. Hunt, $0$-0002-0146-6078; A. Athair, $0$-0001-9282-9577; O. Williams, $0$-0002-5470-5028; B. Polagye, $0$-0002-5156-5007.
@@ Appendix A Influence of Submergence Depth on Performance
As mentioned in Section 2.2, $s/h$ was maximized at each blockage to minimize the impact of ventilation on array performance, particularly at the highest blockage ratios. However, this resulted in variation of $s/h$ across the tested $\beta$ (table 1). To assess the sensitivity of array performance to submergence depth, the array was also tested at a range of $s/h$ at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ through variation of the rotors’ vertical position in the water column. The ADV position was likewise adjusted such that $U_{\infty}$ was always sampled at the turbine midplane.

### v2

@@ V Prediction of Confined Performance
To quantify the effectiveness of this forecasting method, the measured array performance at $\beta_{1}=30.0\%-55.0\%$ (Figure 6) is used to predict array performance at $\beta_{2}$, with the resulting $C_{T,2}$ and $C_{P,2}$ at each operating point compared to the experimentally measured performance (denoted as $C_{T,0}$ and $C_{P,0}$). The predicted thrust coefficients at $\beta_{2}=30.0\%$, $45.0\%$ and $55.0\%$ and the corresponding difference between the predicted and measured thrust coefficients are given in Figure 19. The predictions are within $\sim\!0.2$ of measurements at $\beta_{2}=30.0\%$ and $45.0\%$ (Figures 19 and 19) and within $\sim\!0.3$ at $\beta_{2}=55.0\%$ (Figure 19). In general, the magnitude of the thrust prediction error increases with the difference between $\beta_{1}$ and $\beta_{2}$ (i.e., predictions further from the starting blockage are less accurate), and the absolute prediction error tends to grow as $\beta_{2}$ increases for the same difference between $\beta_{1}$ and $\beta_{2}$. In relative terms, these errors are appreciable at lower $\lambda_{2}$ where $C_{T}$ is relatively low, but decline for higher $\lambda_{2}$. Since the foundation of this forecasting method is Maskell-inspired bluff-body scaling (Eqs. 16, 17 and 18), the differences in the predicted $C_{T,2}$ across $\beta_{1}$ are driven by differences in ${C_{T}}^{b}$ across $\beta_{1}$. Thus, the shapes and relative positions of the ${C_{T}}^{b}-\lambda^{b}$ curves across $\beta_{1}$ (Figure 17) dictate the shapes and relative positions of the resulting $C_{T,2}-\lambda_{2}$ curves when scaled to $\beta_{2}$ (Figures 19, 19 and 19). For the current data set, at low to moderate tip-speed ratios this corresponds to underpredicted $C_{T,2}$ if $\beta_{1}>\beta_{2}$ and overpredicted $C_{P,2}$ if $\beta_{1}<\beta_{2}$, except at relatively high $\lambda_{2}$ where this trend inverts.
@@
The corresponding predicted power coefficients and prediction error are shown in Figure 20. The predicted power coefficients are within $\sim\!0.05$ of measurements at $\beta_{2}=30.0\%$ (Figure 20), $\sim\!0.1$ at $45.0\%$ (Figure 20), and $\sim\!0.2$ at $\beta_{2}=55.0\%$ (Figure 20). The greatest prediction error tends to occur near the performance peak and at tip-speed ratios well beyond the optimal operating point. However, the tip-speed ratio corresponding to maximum performance is well-predicted (Figures 20, 20 and 20). Since only the thrust influences the outcome of the bluff-body forecasting procedure, trends in the $C_{P,2}-\lambda_{2}$ curves with $\beta_{1}$ resemble those for the ${C_{T}}^{b}-\lambda^{b}$ curves.
@@
The proposed forecasting method combines linear momentum theory and the Maskell-inspired bluff-body scaling to enable predictions for confined turbine performance from experimental or simulation data. However, since this method utilizes $u_{b}$ as a link between $\beta_{1}$ and $\beta_{2}$, the prediction accuracy depends on whether the bypass velocity is descriptive of array performance at both $\beta_{1}$ and $\beta_{2}$. Consequently, the effectiveness of this forecasting scheme outside of the tested range of $\beta$ is unknown, and the method is likely less accurate when turbine dynamics depart from a bluff-body model. If there exists a transitional condition at which both a Maskell-inspired LMADT scaling (Figure 17) and traditional Glauert-derived LMADT scaling (Figure 17) provide similar results, then a hybrid method that translates between these flow regimes might allow for forecasting performance across a wider ranges of blockages. However, the existence of such a transitional condition is speculative.
VI Conclusions
@@
This research investigates the performance and near-wake flow fields of a cross-flow turbine array at the upper end of practically achievable blockage ratios and demonstrates the efficacy of theoretical frameworks for describing the dynamics of real turbines under these conditions. The performance and near-wake flow field of a two-turbine array were experimentally characterized over a wide range of operating conditions at array blockage ratios from $30\%$ to $55\%$, holding other non-dimensional flow parameters constant to isolate blockage effects. The observed array performance is consistent with trends observed in prior work while providing a uniquely high-resolution view of the manner in which performance evolves with the blockage ratio at relatively high confinements. This highlights relatively subtle trends, such as linear behavior of both the power and thrust coefficients with blockage for low-to-moderate $\beta$ and $\lambda$. Similarly, the measured near-wake velocity fields quantitatively illustrate how the wake and bypass flows change with $\beta$ and $\lambda$. This work also provides an openly-available set of experimental performance and flow-field data for cross-flow turbines at high blockage, which will aid in the validation of numerical simulations and reduced-order models that seek to examine confinement effects and inform the design of full scale systems that can exploit these physics.
@@
A key outcome of this work is the demonstration that linear momentum actuator disk models can quantitatively describe aspects of real turbine arrays in highly confined flows. It is shown that, at a location $0.6$ turbine diameters downstream of the array, the velocity of the flow bypassing the array measured using PIV is well represented by the bypass velocity predicted from the open-channel LMADT model of Houlsby et al. [1] using the measured array thrust coefficients and inflow conditions as inputs. This result was confirmed for a wide range of blockages and tip-speed ratios, with the bypass velocity calculated from LMADT within approximately $8\%$ of the measured velocity between the turbines at the array centerline at all conditions. The residual deviation between modeled and measured bypass velocities may be a consequence of near-wake mixing that is neglected in LMADT. Following Whelan et al. [2] and inspired by the bluff-body theory of Maskell [3], scaling the array power and thrust coefficients by this analytical bypass velocity collapses these coefficients across the tested blockage and tip-speed ratios. These results highlight similar dynamics between high-blockage turbine arrays and bluff bodies, and demonstrate that, despite their inherent simplicity, LMADT models can effectively model aspects of the flow around non-ideal turbines in confined flows and provide insights into the salient hydrodynamics that govern turbine performance. Leveraging the connection between array performance and the bypass velocity, an analytical forecasting method based on LMADT is presented by which known turbine performance at one confinement ($\beta_{1}$) can be used to predict turbine performance at a different confinement ($\beta_{2}$) with moderate accuracy. This forecasting method provides a pathway for estimating turbine performance under confined conditions based on the dominant dynamics of high-blockage arrays.
@@
Although this work focuses on cross-flow turbines, the characteristic fluid dynamics of confinement are not exclusive to this archetype, and the analytical models employed are agnostic to turbine geometry. However, the design of the rotor is expected to influence how these characteristic dynamics change with blockage and the tip-speed ratio [13], and the effectiveness of Maskell-inspired bluff-body scaling and forecasting are likely dependent on turbine geometry and operating condition. It is recommended that future work explore the broader application of these analytical methods to turbines at lower confinement, as well as to different turbine designs. Additionally, although this work focuses on the time-averaged array performance and near-wake flow fields, the unsteady fluid dynamics of cross-flow turbines also inform their performance and the properties of the near-wake. Future research should examine how the phase-resolved performance of cross-flow turbines develops with confinement, and how these dynamics influence the temporal and spatial evolution of the flow field in the near- and far-wake regions.
Acknowledgements
@@
This work was supported by the United States Advanced Research Projects Agency – Energy (ARPA-E) under award number DE-AR0001441. The authors would like to thank Gregory Talpey, Gemma Calandra, and Corey Crisp for their assistance in commissioning the high-confinement test-rig, as well as help with data collection. Abigale Snortland, Hannah Ross, Jennifer Franck, and Mukul Dave are acknowledged for insightful conversations regarding the results in this work. We are grateful to the Alice C. Tyler Charitable Trust for upgrades to the experimental flume, including closed loop heating and cooling.
Data Availability
@@
The data that support the findings of this study are openly available in the Dryad data repository at https://doi.org/10.5061/dryad.dfn2z35b5 [72] (DOI pending activation upon acceptance). The MATLAB code used to implement the models of Houlsby et al. [1] and Steiros et al. [24] is available on GitHub at https://github.com/aidan-hunt/turbine-confinement-models.
Supplementary Material
@@
See supplementary material at [URL will be inserted by publisher] for lateral force measurements, estimates of hydraulic efficiency, application of Steiros et al.’s model to the experimental data, a visualization of the solution space for the bluff-body forecasting method, and additional figures of PIV flow fields at $0.6D$ and $1.5D$ downstream of the array.
Author contributions (CRediT)
@@
A. Hunt: Conceptualization, Investigation, Validation, Methodology, Formal Analysis, Visualization, Software, Writing - Original Draft and Review & Editing. A. Athair: Investigation, Validation, Methodology, Formal Analysis, Visualization, Software, Writing - Original Draft and Review & Editing. O. Williams: Conceptualization, Methodology, Supervision, Writing - Review & Editing B. Polagye: Conceptualization, Methodology, Supervision, Writing - Original Draft and Review & Editing, Resources, Funding Acquisition.
@@ Appendix A Influence of Submergence Depth on Performance
As mentioned in Section II.2, $s/h$ was maximized at each blockage to minimize the impact of ventilation on array performance, particularly at the highest blockage ratios. However, this resulted in variation of $s/h$ across the tested $\beta$ (Table 1). To assess the sensitivity of array performance to submergence depth, the array was also tested at a range of $s/h$ at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ through variation of the rotors’ vertical position in the water column. The ADV position was likewise adjusted such that $U_{\infty}$ was always sampled at the turbine midplane.